OpenStudy (anonymous):

can anyone Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 5x2 – 100x + 455 = 0

OpenStudy (anonymous):

5(x^2-20x+91)=0 5[(x-10)^2-100+91]=0 5[(x-10)^2-9]=0 5(x-10)^2-45=0 (x-10)^2=9